Mphar34576:

I’m order for the data in the table to represent a linear function with a rate of +5 what must be the value of a

Mphar34576:

1 attachment
surjithayer:

a-13=13-a find a

mhanifa:

This would be: a - 13 = 5 or 23 - a = 5